A store sells customizable backpacks. The customer can choose a size (S, M, L ), then a pattern (solid, striped or checkered), then the number of pockets (1,2,3,4), and finally they ask you to choose one of 5 different colors. How many possible backpacks can be made if the only restriction is that small backpacks must have 1 or 2 pockets?
Answer #1

No. of possible backpacks of small size (S) = 3*2*5 = 30

( 3 possible patterns, 2 possible no. of pockets, 5 possible colors)

No. of possible backpacks of size M and L = 3*4*5 = 60 each

Thus, total number of possible backpacks = 30 + 60 + 60 = 150